# Yue Liao  Pengfei Zhou  Siyuan Huang  Donglin Yang  Shengcong Chen Yuxin Jiang  Yue Hu Jingbin Cai  Si Liu  Jianlan Luo  Liliang Chen Shuicheng Yan  Maoqing Yao  Guanghui Ren†⋄

### v1

@@ Introduction
Embodied agents that sense, reason, and act in the physical world represent the next frontier of AI systems. At its core, a fundamental research challenge remains: developing scalable and robust robotic manipulation capabilities - the ability to purposefully interact with and control the physical environment through selective contact (Mason, 2001). Despite considerable progress has been made in this domain, ranging from analytic (Berenson et al., 2009; Stilman, 2007), model-based frameworks (Ebert et al., 2018; Janner et al., 2019; Nagabandi et al., 2020) to data-driven approaches that learn manipulation policies from large-scale datasets (Kim et al., 2024; Black et al., 2024; Brohan et al., 2023; Bu et al., 2025b), existing systems typically rely on a patchwork of separate data-collection, training, and evaluation stages. Each stage demands bespoke infrastructure, manual curation, and task-specific tuning; the resulting friction could potentially slow down iteration, obscure failure modes, and impede reproducibility at scale. These fragmented stages underscore the absence of an integrated framework capable of learning and evaluating manipulation policies in a unified manner.
@@
To this end, we introduce Genie Envisioner (GE), a unified platform that collapses robot sensing, policy learning, and evaluation into a single closed-loop video generative world model, as illustrated in Figure 1. At its core lies GE-Base, an instruction-conditioned, multi-view video diffusion model trained on approximately 3,000 hours of video-language paired data spanning over one million real-world robotic manipulation episodes from the AgiBot-World-Beta dataset (Bu et al., 2025a). Conditioned on robot’s visual observations, GE-Base autoregressively generates video chunks that capture the temporal evolution of manipulation behaviors following high-level instructions. Leveraging robotic domain adaptation pretraining, GE-Base establishes a mapping from language instructions to an embodied visual space, capturing the essence of robotic manipulation by modeling the spatial, temporal, and semantic regularities of real-world interactions. It achieves this by inferring latent trajectories that jointly encode the robot’s perceptual inputs and the anticipated evolution of the scene under plausible action sequences. To bridge the gap between visual representations and executable robotic control, we introduce GE-Act, a lightweight parallel flow-matching action model. GE-Act translates visual latent features, conditioned on language instructions, into fine-grained and low-latency motor commands, enabling direct and efficient mapping from perception and instruction to executable physical actions. Beyond policy learning, simulation plays a critical role in enabling scalable training, safety validation, and fast iteration for robotic systems. To this end, we introduce GE-Sim, which leverages the embodied video generation capabilities of GE-Base and repurposes its generative dynamics into an action-conditioned world simulator. GE-Sim supports closed-loop policy evaluation through video-based simulation, achieving speeds significantly faster than real-world execution. After designing the core foundation model, a critical challenge remains: evaluating whether the generated videos faithfully simulate robotic behaviors. This requires moving beyond generic perceptual metrics to assess whether the synthesized behaviors are both physically grounded and semantically aligned with the given instructions. To address this, we propose the Embodied World Model Benchmark (EWMBench), a principled evaluation suite that directly benchmarks video generative neural world simulators in terms of visual fidelity, physical consistency, and instruction–action alignment. Therefore, GE constructs a unified video-based robotic vision space that facilitates the learning, simulation, and evaluation of action policies within a perceptually grounded framework. Different from mainstream vision-language-action (VLA) methods (Kim et al., 2024; Black et al., 2024) that rely on vision-language models (VLMs) (Bai et al., 2025; Chen et al., 2024; Abouelenin et al., 2025) to map visual inputs into a semantic linguistic space and learn action policies from this language-centric representation, GE constructs a vision-centric space through generative video modeling. This space preserves detailed spatial and temporal cues, enabling more faithful modeling of robot–environment dynamics and supporting end-to-end policy learning and evaluation within a single, coherent platform.
@@
Together, these contributions position Genie Envisioner as a practical, scalable foundation for real-world manipulation, facilitating downstream research. All code, pretrained models, and the complete EWMBench suite will be open-sourced upon publication to accelerate future research.
@@
In this section, we present GE-Base, the core component of Genie-Envisioner. Our objective is to extend the predictive capacity of general video generation models toward constructing an embodied predictive representation—a unified generative formulation that anticipates future robot-environment interactions conditioned on task instructions and grounded in the agent’s physical embodiment. To this end, we formulate robotic video world modeling as a text-and-image-to-video generation problem: given a language instruction and an initial visual observation, the model forecasts future video segments that reflect plausible and coherent robotic behaviors. A key design feature of GE-Base is its sparse memory mechanism, which augments the current visual input with long-term historical context, enabling stronger temporal reasoning through a unified visual condition. Built upon this formulation, GE-Base adopts a video diffusion transformer architecture and incorporates a robotic-adaptive pretraining strategy that transfers knowledge from generic video datasets into the embodied robotic domain. We demonstrate the effectiveness of GE-Base on real-world robotic manipulation video generation. Experimental results show that GE-Base produces instruction-aligned, temporally coherent video sequences that generalize well across diverse manipulation tasks and embodiments.
@@
Beyond validating GE on the in-domain AgiBot G1 platform, we assess GE’s ability to generalize across different embodiments, an essential step toward developing a versatile robotic foundation model. Specifically, we evaluate GE on two widely used robotic platforms in manipulation research: the Franka arm and the Agilex Cobot Magic system. To maintain consistency with our dual-arm framework, both platforms are configured accordingly.
@@
Figure 11: Performance comparison of two deformable object folding tasks on Agilex Cobot Magic.

### v2

@@ Introduction
Embodied agents that sense, reason, and act in the physical world represent the next frontier of AI systems. At its core, a fundamental research challenge remains: developing scalable and robust robotic manipulation capabilities - the ability to purposefully interact with and control the physical environment through selective contact (Mason, 2001). Despite considerable progress has been made in this domain, ranging from analytic (Berenson et al., 2009; Stilman, 2007), model-based frameworks (Ebert et al., 2018; Janner et al., 2019; Nagabandi et al., 2020) to data-driven approaches that learn manipulation policies from large-scale datasets (Kim et al., 2024; Black et al., 2024; Brohan et al., 2023; Bu et al., 2025b), existing systems typically rely on a patchwork of separate data‑collection, training, and evaluation stages. Each stage demands bespoke infrastructure, manual curation, and task‑specific tuning; the resulting friction could potentially slow down iteration, obscure failure modes, and impede reproducibility at scale. These fragmented stages underscore the absence of an integrated framework capable of learning and evaluating manipulation policies in a unified manner.
@@
To this end, we introduce Genie Envisioner (GE), a unified platform that collapses robot sensing, policy learning, and evaluation into a single closed-loop video generative world model, as illustrated in Figure 1. At its core lies GE‑Base, an instruction-conditioned, multi-view video diffusion model trained on approximately 3,000 hours of video-language paired data spanning over one million real-world robotic manipulation episodes from the AgiBot‑World‑Beta dataset (Bu et al., 2025a). Conditioned on robot’s visual observations, GE‑Base autoregressively generates video chunks that capture the temporal evolution of manipulation behaviors following high-level instructions. Leveraging robotic domain adaptation pretraining, GE‑Base establishes a mapping from language instructions to an embodied visual space, capturing the essence of robotic manipulation by modeling the spatial, temporal, and semantic regularities of real-world interactions. It achieves this by inferring latent trajectories that jointly encode the robot’s perceptual inputs and the anticipated evolution of the scene under plausible action sequences. To bridge the gap between visual representations and executable robotic control, we introduce GE‑Act, a lightweight parallel flow-matching action model. GE‑Act translates visual latent features, conditioned on language instructions, into fine-grained and low-latency motor commands, enabling direct and efficient mapping from perception and instruction to executable physical actions. Beyond policy learning, simulation plays a critical role in enabling scalable training, safety validation, and fast iteration for robotic systems. To this end, we introduce GE-Sim, which leverages the embodied video generation capabilities of GE‑Base and repurposes its generative dynamics into an action-conditioned world simulator. GE-Sim supports closed-loop policy evaluation through video-based simulation, achieving speeds significantly faster than real-world execution. After designing the core foundation model, a critical challenge remains: evaluating whether the generated videos faithfully simulate robotic behaviors. This requires moving beyond generic perceptual metrics to assess whether the synthesized behaviors are both physically grounded and semantically aligned with the given instructions. To address this, we propose the Embodied World Model Benchmark (EWMBench), a principled evaluation suite that directly benchmarks video generative neural world simulators in terms of visual fidelity, physical consistency, and instruction–action alignment. Therefore, GE constructs a unified video-based robotic vision space that facilitates the learning, simulation, and evaluation of action policies within a perceptually grounded framework. Different from mainstream vision-language-action (VLA) methods (Kim et al., 2024; Black et al., 2024) that rely on vision-language models (VLMs) (Bai et al., 2025; Chen et al., 2024; Abouelenin et al., 2025) to map visual inputs into a semantic linguistic space and learn action policies from this language-centric representation, GE constructs a vision-centric space through generative video modeling. This space preserves detailed spatial and temporal cues, enabling more faithful modeling of robot–environment dynamics and supporting end-to-end policy learning and evaluation within a single, coherent platform.
@@
Together, these contributions position Genie Envisioner as a practical, scalable foundation for real‑world manipulation, facilitating downstream research. All code, pretrained models, and the complete EWMBench suite will be open‑sourced upon publication to accelerate future research.
@@
In this section, we present GE-Base, the core component of Genie-Envisioner. Our objective is to extend the predictive capacity of general video generation models toward constructing an embodied predictive representation—a unified generative formulation that anticipates future robot-environment interactions conditioned on task instructions and grounded in the agent’s physical embodiment. To this end, we formulate robotic video world modeling as a text-and-image-to-video generation problem: given a language instruction and an initial visual observation, the model forecasts future video segments that reflect plausible and coherent robotic behaviors. A key design feature of GE‑Base is its sparse memory mechanism, which augments the current visual input with long-term historical context, enabling stronger temporal reasoning through a unified visual condition. Built upon this formulation, GE‑Base adopts a video diffusion transformer architecture and incorporates a robotic-adaptive pretraining strategy that transfers knowledge from generic video datasets into the embodied robotic domain. We demonstrate the effectiveness of GE‑Base on real-world robotic manipulation video generation. Experimental results show that GE‑Base produces instruction-aligned, temporally coherent video sequences that generalize well across diverse manipulation tasks and embodiments.
@@
Analysis. To systematically analyze our GE model, we conduct real-world robotic manipulation experiments on the AgiBot-G1 platform. We select a stable and controllable task, "grasping a red cylinder from the table and placing it into a paper cup with fixed positions", using a dataset of 305 demonstrations. All models are trained for 40,000 steps under the same protocol. Our analysis focuses on the role of pretraining in action policy prediction, comparing general video pretraining with in-domain embodied pretraining (AgiBot-World-Beta). As shown in Table 3.4, training from scratch or adapting from a general video model such as LTX-Video yields near-zero success.
Table 1: Analysis of Pre-training. ‘S’ denotes inclusion of robot state; ‘VidAW’ indicates initialization from GE-Base, ‘VidAda’ indicates task-specific video adaptation.
E2E SR
VidAW VidAda w/ $\mathcal{S}$ w/o $\mathcal{S}$ w/ $\mathcal{S}$ w/o $\mathcal{S}$
0.15 0.30 0.05 0.11
0 0.05 0 0
0.81 0.49 0.64 0.26
0.89 0.37 0.76 0.37
In contrast, in-domain pretraining achieves 64 SR and 81 E2E, which further improve to 76% and 89% when combined with general video pretraining. We further validate the effectiveness of incorporating robot state as input, which yields additional performance gains. However, when applied directly to general video-pretrained models, the inclusion of state information reduces performance due to short-cut learning effects. These results demonstrate that the GE-Base pretrained world model offers strong representations and serves as a solid foundation for action policy prediction.
@@
Beyond validating GE on the in-domain AgiBot G1 platform, we assess GE’s ability to generalize across different embodiments, an essential step toward developing a versatile robotic foundation model. Specifically, we evaluate GE on two widely used robotic platforms in manipulation research: the Franka arm and the Agilex Cobot Magic system, as well as a dual-arm simulator, RoboTwin. To maintain consistency with our dual-arm framework, all platforms are configured accordingly.. To maintain consistency with our dual-arm framework, both platforms are configured accordingly.
@@
Figure 11: Comparison of Task-Specific Manipulation Performance on Various Embodiments.
@@
Generalization to RoboTwin
We further evaluate the cross-embodiment generalization on the dual-arm simulator RoboTwin (Chen et al., 2025). We adopt an all-in-one strategy, jointly fine-tuning GE-Act on four tasks using 200 demonstrations (50 per task), and directly evaluating this unified model across all tasks. In contrast, baseline methods (Black et al., 2024; Bu et al., 2025a) perform task-specific adaptation. As shown in Figure 11, GE-Act achieves better performance than $\pi_{0}$ and GO-1 on three of the four tasks, despite not using a one-task-one-model setting, and is only slightly behind VLA methods on lift pot. This minor gap may be attributed to task interference introduced by joint training.